In error and measurement if there is an error in diametre then there is same error in radius or different

Grade:12

1 Answers

Arun
6 years ago
As we know
d = 2 r
Taking log both the sides
Log d = Log( 2r)
log d = log 2 + log r
On differentiation
d D/d = 0 + d r/r
Hence 
Percentage change is same